A sample of helium gas occupies a volume of 8.89 L at 59.0°C and 1.26 atm.


If it is desired to increase the volume of the gas sample to 10.5 L, while increasing its pressure to 1.54 atm, the temperature of the gas sample at the new volume and pressure must be 

°C.


1
Expert's answer
2021-10-05T23:58:30-0400

For calculations, temperatures must be converted to Kelvin scale first:

T1 = 59.0 + 273 = 332 K

According to the combined gas law,

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}" . Therefore,


"T_2=\frac{T_1P_2V_2}{P_1V_1}=\frac{332K\times1.54atm\times10.5L}{1.26atm\times8.89L}=479K=479-273=206\degree{C}"


Answer: 206oC